My question is: Marys father is 4 times as old as Mary and 5 years ago he was 7 times as old. How old is each now?
Answer by Mathtut(3670)

let m and f be ages of mary and mary's father
:
f=4m..........eq 1
f-5=7(m-5)....eq 2
:
take f's value from eq 1 and plug it into eq 2
:
4m-5=7m-35
:
3m=30
:
m=10
:
f=4m=4(10)=40
: